OpenStudy (anonymous):

In 1985, college graduates averaged $508 in weekly earnings. This amount has increased by approximately $25 in weekly earnings per year. By contrast, in 1985, people with less than four years of high school averaged $270 in weekly earnings. This amount has only increased by approximately $4 in weekly earnings per year. How many years after 1985 will college graduates be earning three times the amount per week that people with less than four years of high school earn? (Round to the nearest whole number.) In which year will this occur? What will be the weekly earnings for each group at

OpenStudy (anonymous):

Ok, so we need to create two base equations for each the college graduates and high school students with less than four years of school. 1) 508 + 25 x 2) 270 + 4x x representing the number of years so we want three times the amount of the college graduates than high school students 3(508 + 25x) = 270 + 4x 75x + 1524 = 270 + 4x 75x - 4x = -1254 71x = - 1254 x = -1254/ 71 (lets take the negative out and just leave it positive) x = 17.66 x = 18 to check we sub x = 18 back into the equation 75(18) + 1524 = 950 270 + 4(18) =342 958/342 = approx 2.8, this is pretty close to the times 3 greater mark so I believe the answer of $950 for graduates and $342 for undergraduates will be correct.